OpenStudy (anonymous):

Running with an initial velocity of 10m/s, a horse has an average acceleration of -1.85m/s2. How long does it take for the horse to decrease its velocity to 6.9m/s ?

Parth (parthkohli):

Use the following equation of motion:\[v_i + at = v_f\]Where the symbols are their usual meanings: \(v_i\) is the initial velocity. \(v_f\) is the final velocity. \(a\) is the average acceleration. \(t\) is the time.

OpenStudy (anonymous):

So would I just insert all the given variables and get t by itself in order to find the time?? My physics prof. hasn't been super clear in his explaining so he leaves it up to us to figure out.

OpenStudy (phi):

yes, fill in the info. what equation do you get ?

Parth (parthkohli):

Follow phi. You know three of those quantities. You thus have one unknown quantity, \(t\).

OpenStudy (anonymous):

(10m/s) + (-1.85m/s^2) t = 6.9m/s would be the equation

Parth (parthkohli):

That's right. The unit of \(t\) is seconds, so you cancel that sec out with the sec in the denominator.

OpenStudy (anonymous):

So then to the 10 m/s would get added to the 6.9m/s and then divide by -1.85??

OpenStudy (phi):

add -10 to both sides

OpenStudy (phi):

(10m/s) -10 m/s+ (-1.85m/s^2) t = 6.9m/s -10 m/s

OpenStudy (anonymous):

Then I would add 1.85 to both sides correct??

OpenStudy (phi):

no, you said it correctly the first time: and then divide by -1.85??

OpenStudy (anonymous):

1.68 sound right?